OpenStudy (anonymous):

9y√(3) and 9√(3y) are like radicals ? √9= is a raditional #? √-1 is not a real #? The Pythagorean Theorem only refers to right triangles.

OpenStudy (anonymous):

well the sqrt of 9 is three you cannot square root negative numbers without using an imagenary number so it is not a real number

OpenStudy (anonymous):

is the first one like radicals?

OpenStudy (anonymous):

i do not believe so, because the y in one is being square rooted and in the other it is not.

OpenStudy (anonymous):

3√(5) is the symbol of the square root of 5 √ab=√a*√b? is this correct ?

OpenStudy (anonymous):

√a-b = √a-√b

OpenStudy (anonymous):

are these correct ?

OpenStudy (anonymous):

i am pretty positive they are but you might want to get a second opinion.

OpenStudy (anonymous):

but 9 is a rational #?

OpenStudy (anonymous):

yes because nine can be sqrted and you get 3.